OpenStudy (anonymous):

The weekly cost (C) of growing and selling x acres of flowers is approximated by C = 0.25x2 10x + 305. How many acres of flowers would minimize these costs? 20 acres 25 acres 30 acres 40 acres

OpenStudy (anonymous):

@SrBenjaminIsrael @stabar

OpenStudy (anonymous):

I don't even know what that is asking but @Mertsj will...sorry :/

OpenStudy (anonymous):

oh:/

OpenStudy (anonymous):

@ZeHanz is also realy good

OpenStudy (anonymous):

@Anders96

OpenStudy (anonymous):

@jes230

OpenStudy (anonymous):

@Mikemartin

OpenStudy (anonymous):

@Mikemartin

OpenStudy (anonymous):

C = 0.25x^2 - 10x + 305 C' = 0.5x - 10 Set that to zero: 0.5x - 10 = 0 0.5x = 10 x = 10 / 0.5 x = 20

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

Ya

OpenStudy (anonymous):

I had the same problem in school THX